Factor completely
2a^2-5a-3

Found 2 solutions by estron7, addingup:
Answer by estron7(44)

Factor by grouping
(a−3)(2a+1)
Hope it helps

Answer by addingup(3677)

Factor completely: 2a^2-5a-3
1) The coefficient of a^2 is 2 and the constant term is -3.
2) The product of 2 and -3 is -6.
3) The factors of -6 which sum to -5 are 1 and -6.
Let's put it all together:
2a^2-5a-3 = 2a^2-6a+a-3 = a(2a+1)-3(2a+1)
Now we can factor 2a+1 and we get:(2a+1)(a-3)